OpenStudy (anonymous):

The sum to n terms of the series \[\sqrt{2}+\sqrt{8}+\sqrt{18}+\sqrt{32}+......\]

OpenStudy (anonymous):

\sqrt{2} *(n*(n+1)/2)

OpenStudy (anonymous):

Take root 2 common

OpenStudy (anonymous):

So, it becomes root2(1+2+3+4+...)

hartnn (hartnn):

lol! so simple, i can't believe you couldn't do this.....

OpenStudy (anonymous):

sqrt(2)+sqrt(2*4)+sqrt(2*9)+sqrt(2*16)+........ =sqrt(2)+2*sqrt(2)+3*sqrt(2)+4*sqrt(2)+...... =sqrt(2)*{1+2+3+4+....} =sqrt(2)*{n*(n+1)/2}